OpenStudy (anonymous):

Find the dimensions of a rectangle with perimeter 116 m whose area is as large as possible.

OpenStudy (anonymous):

make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square make a square

myininaya (myininaya):

2x+2y=116 where x is length and y is width A=xy solve for x or y from 2x+2y=116 doesn't matter which go ahead and divide both sides by 2 before doing so x+y=58 so we have y=58-x now A=x(58-x)=58x-x^2 A'=58-2x 58-2x=0 2x=58 x=58/2=29 so y=29

myininaya (myininaya):

calculus or algebra??

OpenStudy (anonymous):

Calc

myininaya (myininaya):

ok cool do you understand?

OpenStudy (anonymous):

Yes, myininaya. Thank you so much, very good explanation! Medal earned without question.

OpenStudy (anonymous):

\[P=2l+2w\] \[\frac{P}{2}=l+2\] \[\frac{P}{2}-l=w\] \[A=l\times w\] \[A(l)=l(\frac{P}{2}-l)\] \[A(l)=\frac{Pl}{2}-l^2\]

OpenStudy (anonymous):

a parabola that faced down, vertex is \[\frac{-b}{2a}=\frac{P}{4}\] and so you see you have a square

OpenStudy (anonymous):

btw of course myininaya is correct. however it is entirely unnecessary to use calc to find the vertex of a parabola.